OpenStudy (anonymous):

x^5-x^3-2x^2 is it increasing or decreasing at -2?

OpenStudy (anonymous):

are you saying if x is -2?

OpenStudy (anonymous):

2x^2 is the same as saying x^4, right?

OpenStudy (anonymous):

-2x^2 would be a positive number, if that is what you are asking...

OpenStudy (anonymous):

is someone not supposed to differentiate?

OpenStudy (anonymous):

i would say it's increasing at -2 if i understand the original question...

OpenStudy (anonymous):

thats was my thnking too

OpenStudy (anonymous):

@basketmath: No, its not. 2 * 2^2 = 8, while 2^4 = 16. @mathfrick lolo, Best way to do this is to differentiate. Differentiation of this is 5x^4-3x^2-4x. At x = -2, it is (5*(-2)^4 - 3* (-2)^2 - 4(-2)) = 76 which is positive. so the function is increasing.

OpenStudy (anonymous):

yeah. my thinking was flawed on the 2x^2 equaling x^4. i am trying to focus on too many things. lol. test in a few hours.... i knew it was increasing. my apologies for any confusion.

OpenStudy (anonymous):

i knew differentiation was part of it thanx Tai